# Mutual-Taught for Co-adapting Policy and Reward Models

### v1

@@ Abstract
During the preference optimization of large language models (LLMs), distribution shifts may arise between newly generated model samples and the data used to train the reward model (RM). This shift reduces the efficacy of the RM, which in turn negatively impacts the performance of the policy model (PM). To address this challenge, we propose Mutual-Taught, a self-training method that iteratively improves both the PM and RM without requiring additional human annotation. Our approach mirrors the expectation-maximization (EM) algorithm. In the E-step, the PM is updated using feedback from the current RM, guiding the PM toward a better approximation of the latent optimal preference distribution. In the M-step, we update the RM by constructing training data from the outputs of the PM before and after the E-step update. This process ensures that the RM adapts to the evolving policy distribution. Experimental results demonstrate that this iterative approach leads to consistent improvements in both models. Specifically, our 8B policy model, LLaMA-3-8B-Instruct-MT, achieves a length-controlled win rate of 54.1% on AlpacaEval-2, while our 8B reward model, FsfairX-LLaMA3-RM-MT, performs on par with GPT-4o-2024-08-06 on RewardBench.
@@
Tianyuan Shi1, Canbin Huang1, Fanqi Wan1, Longguang Zhong 1, Ziyi Yang1, Weizhou Shen 2, Xiaojun Quan 1,thanks:   Corresponding authors. Ming Yan2, 1School of Computer Science and Engineering, Sun Yat-sen University, China 2Alibaba Group {shity6, wanfq}@mail2.sysu.edu.cn, quanxj3@mail.sysu.edu.cn, ym119608@alibaba-inc.com
@@ 4.2 Objective of Mutual-Taught
$r_{t}=\operatorname*{argmax}_{r}\mathbb{E}_{x\sim\mathcal{D}_{\text{R}}}\left[\log P_{r}(y_{t}\succ y_{t-1}\!\mid\!x)\right].$
@@ 4.3 Two-Stage Stabilization
If no candidate in this iteration demonstrates sufficient improvement ($\max_{k}w^{(k)}_{t}<\tau$), the iteration halts, and the previous model is preserved.
@@
To adaptively filter noisy comparisons, we establish a variance-aware threshold $\epsilon_{t}=\sqrt{\mathbb{V}_{x\sim\mathcal{D}}[r_{t-1}(y_{t-1};x)]}$ that automatically adjusts to the reward model’s uncertainty level Pace et al. (2024). Only pairs satisfying $|\Delta r(x)|\geq\epsilon_{t}$ are considered high-confidence pseudo-labels. Our filtering strategy removes pairs with $\Delta r(x)\leq-\epsilon_{t}$, as they are confidently identified as noisy samples.
@@ 5.3 Further Analysis
Figure 4: The impact of different reward model training data types on the performance of Mutual-Taught. For brevity, policy-comparison data and self-training data are abbreviated as PC and ST, respectively.
@@
Performance of Mutual-Taught with additional iterations
To examine the impact of iterations on Mutual-Taught, we conduct an additional iteration after the main experiment, using the same training data and experimental parameters. To avoid overfitting from repeated training on the same data, the policy and reward models from each iteration are not used as the starting point for the next. Instead, they are used solely to generate higher-quality training data for the next iteration, with the new iteration’s models starting from the base models. The results of this experiment are shown in Figure 5.
Figure 5: The performance of the policy (left) and the reward (right) models across two rounds. Each round includes two policy updates and one reward model update. For brevity, each policy update is abbreviated (e.g., the first update in Round 1 is denoted as R1–U1).
We note that after the second round, both the policy and the reward models show consistent improvements compared to the first round. Notably, the final reward model outperforms GPT-4o-2024-08-06 on RewardBench, demonstrating that Mutual-Taught achieves even better performance with an additional round. More specifically, in the second iteration, both the policy and reward models utilize preference data generated by their respective fine-tuned predecessors. These higher-quality outputs strengthen the foundation for the E-step (policy updates) and M-step (reward model updates) and result in better alignment between the policy and reward models and enhanced results.
@@ Appendix A Baselines
We compare our approach against the following baseline methods. Offline preference optimization methods: For this category, we implement DPO Rafailov et al. (2023), IPO  Azar et al. (2024) and SimPO Meng et al. (2024). Preference pairs are derived from multiple responses generated by the base policy model, with scores provided by the base reward model. Iterative preference optimization methods: For this category, we implement SPPO Wu et al. (2025) and Meta-Rewarding Wu et al. (2024). Since these methods do not update the reward model, we use all three portions of the dataset for policy model training and run three iterations for iterative methods, i.e., SPPO and Meta-Rewarding. To ensure a fair comparison, the sampling settings used in these experiments match those applied in Mutual-Taught.
@@ Appendix B Training Details
We conduct Mutual-Taught between the policy and reward models for two iterations. In each iteration, both models are trained for one epoch using a cosine learning rate schedule with a warmup ratio of 0.1. All experiments are conducted on 8 NVIDIA A100 GPUs. We follow SimPO Meng et al. (2024) to set the policy sampling and training parameters. Specifically, for policy sampling: the temperature is set to 0.8, $M=5$, and top-p to 0.95. For each policy model iteration, we initialize the model from the previous round and generate responses using the current policy. Preference data is then derived using the reward model at the current iteration. The policy model is optimized via DPO with a beta of 0.01, a batch size of 128, a maximum sequence length of 2,048 tokens, and a learning rate of $7\times 10^{-7}$. A checkpoint is saved every 50 steps for subsequent model selection. For model selection, a fixed evaluation set is constructed prior to the start of the iterations by randomly sampling 2,000 prompts from the UltraFeedback dataset. Among the saved checkpoints, the one with the highest win-rate relative to the initial policy of the current iteration is selected to construct the pseudo-labels. The iteration is terminated if the highest win-rate $w_{t}^{k}$ is less than 60%. For data filtering, the margin threshold is set based on the variance of the reward model scores in the current iteration. Parameter search was conducted over multipliers of 1, 2, and 3 times the variance. The best results were obtained with a threshold set at twice the variance.

### v2

@@ Abstract
During the preference optimization of large language models (LLMs), distribution shifts may arise between newly generated model samples and the data used to train the reward model (RM). This shift reduces the efficacy of the RM, which in turn negatively impacts the performance of the policy model (PM). To address this challenge, we propose Mutual-Taught, a self-training method that iteratively improves both the PM and RM without requiring additional human annotation. Our approach mirrors the expectation-maximization (EM) algorithm. In the E-step, the PM is updated using feedback from the current RM, guiding the PM toward a better approximation of the latent optimal preference distribution. In the M-step, we update the RM by constructing training data from the outputs of the PM before and after the E-step update. This process ensures that the RM adapts to the evolving policy distribution. Experimental results demonstrate that this iterative approach leads to consistent improvements in both models. Specifically, our 8B policy model, Llama-3-8B-Instruct-MT, achieves a length-controlled win rate of 54.1% on AlpacaEval-2, while our 8B reward model, FsfairX-Llama3-RM-MT, performs on par with GPT-4o-2024-08-06 on RewardBench. Our code is available at https://github.com/Stycoo/Mutual-Taught.
@@
Tianyuan Shi1, Canbin Huang1, Fanqi Wan1, Longguang Zhong1, Ziyi Yang1 Weizhou Shen2, Xiaojun Quan1thanks:   Corresponding authors., Ming Yan2 1School of Computer Science and Engineering, Sun Yat-sen University, China 2Alibaba Group {shity6, huangcb3, wanfq, zhonglg5, yangzy39}@mail2.sysu.edu.cn quanxj3@mail.sysu.edu.cn, {shenweizhou.swz, ym119608}@alibaba-inc.com
@@ 4.2 Objective of Mutual-Taught
$r_{t}=\operatorname*{argmax}_{r}\;\mathbb{E}_{x\sim\mathcal{D}_{\text{R}}}\left[\log P_{r}(y_{t}\succ y_{t-1}\mid x)\right]$
@@ 4.3 Two-Stage Stabilization
If no candidate in this iteration demonstrates sufficient improvement ($\max_{k}w^{k}_{t}<\tau$), the iteration halts, and the previous model is preserved.
@@
To adaptively filter noisy comparisons, we establish a variance-aware threshold $\epsilon_{t}=\sqrt{\mathbb{V}_{x\sim\mathcal{D}_{R}}[r_{t-1}(y_{t-1};x)]}$ that automatically adjusts to the reward model’s uncertainty level Pace et al. (2024). Only pairs satisfying $|\Delta r(x)|\geq\epsilon_{t}$ are considered high-confidence pseudo-labels. Our filtering strategy removes pairs with $\Delta r(x)\leq-\epsilon_{t}$, as they are confidently identified as noisy samples.
@@ 5.3 Further Analysis
Figure 4: Impact of different reward model training data types on the performance of Mutual-Taught. For brevity, policy-comparison data and self-training data are abbreviated as PC and ST, respectively.
@@
Performance of Mutual-Taught with additional rounds
To investigate the effect of extending Mutual-Taught beyond the main experimental setup, we conduct an additional round of training using the same dataset and hyperparameters. Each round consists of two policy model updates and one reward model update. Crucially, to mitigate overfitting from repeated training on the same data, the policy and reward models from the previous round are not directly fine-tuned further. Instead, they are used solely to generate higher-quality training data for the next iteration, with the new iteration’s models starting from the base models.
Figure 5: Performance of the policy (left) and the reward (right) models across two rounds. Each round includes two policy updates and one reward model update. For brevity, each policy update is abbreviated (e.g., the first update in Round 1 is denoted as R1–U1).
As shown in Figure 5, both the policy and reward models continue to improve in the second round relative to the first. Notably, the final reward model outperforms GPT-4o-2024-08-06 on RewardBench, demonstrating that Mutual-Taught achieves even better performance with an additional round. More specifically, in the second iteration, both the policy and reward models utilize preference data generated by their respective fine-tuned predecessors. These higher-quality outputs strengthen the foundation for the E-step (policy updates) and M-step (reward model updates) and result in better alignment between the policy and reward models and enhanced results.
@@ 6 Conclusion
Acknowledgements
This work was supported by the National Natural Science Foundation of China (No. 62176270) and the Guangdong Basic and Applied Basic Research Foundation (No. 2023A1515012832).
@@ Appendix A Baselines
We compare our approach against the following baseline methods. Offline preference optimization methods: For this category, we implement DPO Rafailov et al. (2023), IPO  Azar et al. (2024) and SimPO Meng et al. (2024). Preference pairs are derived from multiple responses generated by the base policy model, with scores provided by the base reward model. Iterative preference optimization methods: For this category, we implement iterative DPO Xu et al. (2023), SPPO Wu et al. (2025) and Meta-Rewarding Wu et al. (2024). Since these methods do not update the reward model, we use all three portions of the dataset for policy model training and run three iterations for iterative methods, i.e., SPPO and Meta-Rewarding. To ensure a fair comparison, the sampling settings used in these experiments match those applied in Mutual-Taught.
@@ Appendix B Training Details
We conduct Mutual-Taught between the policy and reward models for two iterations. In each iteration, both models are trained for one epoch using a cosine learning rate schedule with a warmup ratio of 0.1. All experiments are conducted on 8 NVIDIA A100 GPUs. We follow SimPO Meng et al. (2024) to set the policy sampling and training parameters. Specifically, for policy sampling: the temperature is set to 0.8, $M=5$, and top-p to 0.95. For each policy model iteration, we initialize the model from the previous round and generate responses using the current policy. Preference data is then derived using the reward model at the current iteration. The policy model is optimized via DPO with a beta of 0.01, a batch size of 128, a maximum sequence length of 2,048 tokens, and a learning rate of $7\times 10^{-7}$. A checkpoint is saved every 50 steps for subsequent model selection. For model selection, a fixed evaluation set is constructed prior to the start of the iterations by randomly sampling 2,000 prompts from the UltraFeedback dataset. Among the saved checkpoints, the one with the highest win-rate relative to the initial policy of the current iteration is selected to construct the pseudo-labels. The iteration is terminated if the highest win-rate $w_{t}^{k}$ is less than 60%. For data filtering, the margin threshold is set based on the standard deviation of the reward model scores in the current iteration.
@@ Appendix F Pseudo-Label Filtering Methods
Appendix G Evaluation on Additional Benchmarks
To further assess the effectiveness of Mutual-Taught across diverse downstream tasks and evaluation metrics, we conducted additional experiments on four benchmarks from the HuggingFace Open LLM Leaderboard  (Beeching et al., 2023): GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), and TruthfulQA (Lin et al., 2022). The results are summarized in Table 4.
Model GSM8K MMLU HellaSwag TruthfulQA Avg.
Base 75.21 65.71 78.48 51.64 67.76
+ IterDPO 69.71 65.19 80.83 52.91 67.16
+ MT 70.67 64.13 81.37 55.21 67.85
Table 4: Accuracy (%) on additional benchmarks from the HuggingFace Open LLM Leaderboard. Base refers to Llama-3-8B-Instruct; + IterDPO and + MT indicate models fine-tuned with Iterative DPO and Mutual-Taught, respectively.
As shown in Table 4, all preference optimization methods show performance drops on MMLU and GSM8K—likely due to the UltraFeedback dataset’s emphasis on alignment over general knowledge and mathematics. In contrast, there is a consistent improvement on HellaSwag and TruthfulQA. These results suggest that the UltraFeedback dataset is more aligned with tasks requiring commonsense reasoning and truthfulness, and that Mutual-Taught is particularly beneficial in these areas.
Appendix H Threshold Selection for $\tau$ in the E-Step
During early training, the policy model (PM) typically improves markedly after each E-step, reflected by validation win rates well above 50%. In this regime, the updated response $y_{t}$ almost always surpasses its predecessor $y_{t-1}$. Meanwhile, the data-filtering procedure in the M-step discards unreliable preference pairs, keeping the reward model (RM) aligned with the evolving PM distribution. As optimization advances, incremental gains taper off and the win rate converges toward 50%. Distinguishing successive policies then becomes difficult, and marginally noisy pairs may impair the RM. To prevent over-optimization while preserving meaningful updates, we introduce a win-rate threshold $\tau$: Above 50% to prevent over optimization, while not excessively high to ensure continued meaningful optimization.
To further evaluate $\tau$’s effectiveness in triggering timely early stopping during performance degradation, we extended our experiment by dividing the original set of 40,000 prompts (previously used in two iterations) into four subsets and conducting four iterations of PM training. The results are summarized in Table 5.
Iter. MS Win (%) ES AlpacaEval-2 LC (%)
1 63.5 No 34.7
2 67.3 No 41.0
3 65.1 No 44.9
4 57.7 40.3
Table 5: Impact of early-stop threshold $\tau$ across iterations. "MS Win" denotes the win rate (%) of the selected model in model selection (MS), "ES" indicates whether early stopping (ES) was applied during the iteration.
As shown in the Table 5, the PM exhibits performance degradation in the fourth iteration. By setting $\tau=60\%$, based on the win rate (63.5%) from the first PM iteration, early stopping was successfully triggered in the fourth iteration. Consequently, the model from the third EM iteration was selected as the final model. This demonstrates how $\tau$ effectively ensures early stopping at the appropriate point when performance degradation is detected.